An executive invests $28,000, some at 8% and some at 4% annual interest. If he receives an annual return of $1,720, how much is invested at each rate?
Answer by TimothyLamb(4379)

x = dollar amount invested at 8%
y = dollar amount invested at 4%
---
x + y = 28000
0.08x + 0.04y = 1720
---
put the system of linear equations into standard form
---
x + y = 28000
0.08x + 0.04y = 1720
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 15000
y = 13000
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---